• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 23:51:22 on September30,2010

    Entry number 333272

    keyword=Swing shift summary

    We got beam at around 6PM. Here's what we did:

    - Harp scan
    - Lumi tests: beam current scan
    - BPM calibration vs harps
    - Moller magnet alignment (by Sasha and Jixie)
    - Arc harp scanners (by Arun)

    We found that spot++ is not working (we lost ROC4 again), so we could not do raster calibration (postponed probably for tomorrow).

    We'll tune the beam through the chicane later to check Compton rates.